Two thousand tickets are sold  each. One ticket will win , two tickets will win  each and three tickets will win  each. Let  denotes the net gain from the purchase of a randomly selected ticket


1
Expert's answer
2021-03-08T19:06:01-0500

Total price of Tickets =$2000


Probability, "P(X=1)=\dfrac{100}{2000}=0.05"


"P(X=2)=\dfrac{1000}{2000}=0.10"


"P(X=3)=\dfrac{300}{2000}=0.15"


The probability Distribution table is given by-



Net gain = "\sum E(X)=0.05+0.20+0.45=0.70."